OpenStudy (anonymous):

solve 4 (a + 2) = 14 - 2 (3-2 a) no solution all real numbers -2 -1

OpenStudy (anonymous):

will fan and medal

Vocaloid (vocaloid):

start by distributing 4(a+2) = ?

OpenStudy (anonymous):

@Vocaloid thats why i cant seem to do :(

Vocaloid (vocaloid):

hint: A(B+C) = A*B + A*C 4(a+2) = ____?

OpenStudy (anonymous):

6a?

Vocaloid (vocaloid):

follow the pattern

Vocaloid (vocaloid):



OpenStudy (anonymous):

um... but u cant add A and the 2

OpenStudy (anonymous):

wait.. it would be 6+A?

Vocaloid (vocaloid):

nope, not quite...

Vocaloid (vocaloid):



OpenStudy (anonymous):

oh.. 8a or 8+a?

Vocaloid (vocaloid):

let A = 4 let B = a let C = 2 A*B + A*C = ?

OpenStudy (anonymous):

4*a is 4a.. and then 4*2 is 8.. so it would be 8a?

OpenStudy (anonymous):

ugh.. im stupid.. :/

OpenStudy (anonymous):

@Kitty449

OpenStudy (anonymous):

helpz me? D:

Vocaloid (vocaloid):

right, you got 4a and 8, so it's 4a + 8, ready to keep going?

OpenStudy (anonymous):

if u give me the answer.. ill medal u.. pls D:

OpenStudy (anonymous):

SOMEONE!!

OpenStudy (anonymous):

hold on my computer wasnt loading

OpenStudy (anonymous):

all real numbers

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

am i getting these right?

OpenStudy (anonymous):

i think so yes